# An operator splitting analysis of Wasserstein–Fisher–Rao gradient flows

### v1

@@ 6.2 Convergence speed: FR-W
We assume w.l.o.g. that $g(\eta_{\tau})$ is symmetric about 0 and that $g(\eta_{\tau})\geq 0$ since $g(\eta_{\tau})$ is strongly convex and thus lower bounded. Recall that
$S_{\textrm{W}}(t,v)(x)$ $=\int(4\uppi t)^{-d/2}\exp\left(-\frac{|x-y-\int_{0}^{t}\nabla\log\pi(X_{s})ds|^{2}}{4t}\right)v(y)dy,$
where $(X_{t})_{t\geq 0}$ solves $dX_{t}=\nabla\log\pi(X_{t})dt+\sqrt{2}dW_{t}$. It follows that $S_{\textrm{W}}\left(\gamma-\tau,g(\eta_{\gamma})\right)\geq 0$ is also symmetric about 0. Proceeding as in the proof of Lemma 6.1 gives $\text{Cov}_{\eta_{\tau}}\left(S_{\textrm{W}}\left(\gamma-\tau,g(\eta_{\gamma})\right),|\nabla g(\eta_{\tau})|^{2}\right)>0$ from which the result follows.
7 Conclusion & Future Work
Operator splitting methods lie at the heart of many gradient flow based algorithms, in particular, for discretisations of the Wasserstein–Fisher–Rao gradient flow. We examine errors introduced by splitting schemes whereby the Wasserstein and Fisher–Rao operators are solved separately (and exactly). We develop variational formulae for the split schemes and identify a perturbation which modifies the dynamics. It is demonstrated that the order in which these operators are evaluated can have a considerable impact on convergence speed, without affecting $\pi$-invariance. This impact is unsurprisingly more pronounced when the chosen step size is large, but not so large that it is nearly equivalent to evaluating either the Wasserstein or Fisher-Rao flow alone. In practical terms, this means that a sequence of alternating operators with a fixed step size $\gamma$ can be re-arranged (e.g. such that $m$ consecutive steps of size $\gamma$ of the same operator are evaluated) in order to reduce time to convergence, with no extra computational burden.
We focus primarily on characterising this behaviour in the Gaussian case and then more broadly in the strongly log-concave case, showing that although the exact Wasserstein–Fisher–Rao flow preserves log-concavity uniformly in time, under some conditions. Decay rates of the exact WFR flow are then characterised in terms of the symmetrised $\mathrm{KL}$, using the derived log-concavity constants and new functional inequalities for Fisher–Rao flows in [carrillo_fisher-rao_2024].
There are several avenues for future work, including (but not limited to) 1) quantifying the impact of further space and time discretisations of the W and FR flows, which are currently assumed to be evaluated exactly; 2) using the derived results to develop adaptive step size splitting schemes and to determine which ordering is preferrable and finally 3) in more precisely quantifying the acceleration/deceleration of symmetrised $\mathrm{KL}$ due to splitting.
Acknowledgements
The authors would like to thank Josh Bon, Sam Power, Florian Maire, Daniel Paulin, Upanshu Sharma for helpful discussions and in particular Sinho Chewi for pointing to reference [Kolesnikov2001] and Andre Wibisono for directing us to Lemma 16 in [liang_characterizing_2025].
Appendix A Auxiliary lemmas
Lemma A.1.
Let Assumption 1 hold. Recall the $k$th commutator $\mathcal{G}^{k}\eta_{\gamma}:=[f_{\textrm{W}},\mathcal{G}^{k-1}\eta_{\gamma}]$ for $k=1,2,\dots$ with $\mathcal{G}^{0}:=f_{\textrm{FR}}$ and $\mathcal{G}^{1}$ as defined in (2.17). The following identity holds
$\int_{0}^{\gamma}e^{(\gamma-\tau)f_{\textrm{W}}}\mathcal{G}^{1}\eta_{\tau}d\tau=\sum_{k=1}^{\infty}\frac{\gamma^{k}}{k!}\mathcal{G}^{k}(\eta_{\gamma}).$
Proof.
Define $f(\tau)=e^{(\gamma-\tau)f_{W}}\mathcal{G}^{1}\eta_{\tau}$ and consider a Taylor expansion around point $\gamma$ of $f$, $f(\tau)=\sum_{k=0}^{\infty}\frac{f^{(k)}(\gamma)}{k!}(\tau-\gamma)^{k}$. Let us compute the derivatives of $f$. Applying the chain rule we have
@@ Appendix A Auxiliary lemmas
$\partial_{s}f_{\textrm{W}}(f_{\textrm{FR}}(\eta_{s}))=f_{\textrm{W}}^{\prime}(f_{\textrm{FR}}(\eta_{s}))\left[\partial_{s}f_{\textrm{FR}}(\eta_{s})\right]=f_{\textrm{W}}(f^{\prime}_{\textrm{FR}}(\eta_{s})[\partial_{s}\eta_{s}])=f_{\textrm{W}}(f^{\prime}_{\textrm{FR}}(\eta_{s})[f_{\textrm{W}}(\eta_{s})])$ (A.1)
@@
$\partial_{s}f_{\textrm{FR}}^{\prime}(\eta_{s})[f_{\textrm{W}}(\eta_{s})]$ $=f_{\textrm{FR}}^{\prime\prime}(\eta_{s})(\partial_{s}\eta_{s},f_{\textrm{W}}(\eta_{s}))+f_{\textrm{FR}}^{\prime}(\eta_{s})[\partial_{s}f_{\textrm{W}}(\eta_{s})]$ (A.2)
@@
Combining (A.1)–(A.2) it the immediately follows that
@@
$\partial_{s}\mathcal{G}^{1}\eta_{s}-f_{\textrm{W}}(\mathcal{G}^{1})(\eta_{s})$ $=-[f_{\textrm{W}},[f_{\textrm{W}},f_{\textrm{FR}}]](S_{\textrm{W}}(s,u_{\gamma}))=-\mathcal{G}^{2}\eta_{s},$
@@
using the chain rule for operators.
Following the same steps as before we have $\partial_{s}\mathcal{G}^{2}\eta_{s}-f_{\textrm{W}}(\mathcal{G}^{2}\eta_{s})=-\mathcal{G}^{3}\eta_{s}$ and thus $f^{(2)}(s)=e^{(\gamma-s)f_{W}}\mathcal{G}^{3}\eta_{s}$. A recursive argument then gives
@@
$=\gamma\mathcal{G}^{1}\eta_{\gamma}+\sum_{k=1}^{\infty}\frac{\mathcal{G}^{k+1}\eta_{\gamma}}{k!}\frac{\gamma^{k+1}}{k+1}$
$=\sum_{k=1}^{\infty}\frac{\gamma^{k}}{k!}\mathcal{G}^{k}(\eta_{\gamma}).$
@@
Let Assumption 1 hold. Then the $k$th commutator $\mathcal{G}^{k}\eta_{\gamma}:=[f_{\textrm{W}},\mathcal{G}^{k-1}\eta_{\gamma}]$ for $k=1,2,3,\dots$ with $\mathcal{G}^{0}:=f_{\textrm{FR}}$ and $\mathcal{G}^{1}$ as defined in (2.17) has a Fisher-Rao structure,
$\mathcal{G}^{k}\eta_{\gamma}=\eta_{\gamma}\left(g_{k}(\eta_{\gamma})-\mathbb{E}_{\eta_{\gamma}}[g_{k}]\right)$
where
$g_{k}(\eta_{\gamma})$ $=-g_{k-1}^{\prime}f_{\textrm{W}}(\eta_{\gamma})+\nabla\log\frac{\eta_{\gamma}}{\pi}\cdot\nabla g_{k-1}(\eta_{\gamma})+\frac{1}{\eta_{\gamma}}\nabla\cdot(\eta_{\gamma}\nabla g_{k-1}(\eta_{\gamma})),\quad\text{for}\enskip k=1,2,3,\dots$
$g_{0}(\eta_{\gamma})$ $=-\log\frac{\eta_{\gamma}}{\pi}$
and $g_{k-1}^{\prime}f_{\textrm{W}}(\eta)$ denotes the Fréchet derivative of $g_{k-1}$ in the direction $f_{\textrm{W}}(\eta)$.
@@
Recall $g_{0}=-\log\frac{\eta_{\gamma}}{\pi}$, where we drop the dependence on $\eta_{\gamma}$ to simplify notation, so that
$f_{\textrm{FR}}(\eta_{\gamma})=\eta_{\gamma}(g_{0}-\mathbb{E}_{\eta_{\gamma}}[g_{0}])$
We will proceed inductively, starting with $\mathcal{G}^{1}$,
$[f_{\textrm{W}},f_{\textrm{FR}}]\eta_{\gamma}$ $=f_{\textrm{W}}^{\prime}f_{\textrm{FR}}(\eta_{\gamma})-f_{\textrm{FR}}^{\prime}f_{\textrm{W}}(\eta_{\gamma}).$ (A.3)
Since $f_{\textrm{W}}$ is a bounded linear operator, its Fréchet derivative is itself and
$f_{\textrm{W}}^{\prime}f_{\textrm{FR}}(\eta_{\gamma})$ $=\nabla\cdot\left(\nabla\eta_{\gamma}(g_{0}-\mathbb{E}_{\eta_{\gamma}}[g_{0}])\right)-\nabla\cdot\left(\eta_{\gamma}(g_{0}-\mathbb{E}_{\eta_{\gamma}}[g_{0}])\nabla\log\pi\right)$ (A.4)
$=\nabla\cdot(\nabla\eta_{\gamma}g_{0})-\nabla\cdot(\eta_{\gamma}g_{0}\nabla\log\pi)-\mathbb{E}_{\eta_{\gamma}}[g_{0}]f_{\textrm{W}}(\eta_{\gamma})$
@@
$=\nabla g_{0}\cdot\nabla\eta_{\gamma}+\eta_{\gamma}\Delta g_{0}+\left(\nabla\eta_{\gamma}-\eta_{\gamma}\nabla\log\pi\right)\cdot\nabla g_{0}+(g_{0}-\mathbb{E}_{\eta_{\gamma}}[g_{0}])f_{\textrm{W}}(\eta_{\gamma})$
@@
The Fréchet derivative of $f_{\textrm{FR}}$ can be obtained by computing its directional derivative
$\frac{d}{d\epsilon}\left.f_{\textrm{FR}}(\eta_{\gamma}+\epsilon h)\right|_{\epsilon=0}$ $=\frac{d}{d\epsilon}\left.-(\eta_{\gamma}+\epsilon h)\log\frac{\eta_{\gamma}+\epsilon h}{\pi}+(\eta_{\gamma}+\epsilon h)\int\log\frac{\eta_{\gamma}+\epsilon h}{\pi}(\eta_{\gamma}+\epsilon h)\right|_{\epsilon=0}$ (A.5)
@@
So then since $g_{0}^{\prime}h=-\frac{h}{\eta_{\gamma}}$,
@@
$[f_{\textrm{W}},f_{\textrm{FR}}]\eta_{\gamma}$ $=(g_{0}-\mathbb{E}_{\eta_{\gamma}}[g_{0}])f_{\textrm{W}}(\eta_{\gamma})+\nabla\cdot(\eta_{\gamma}\nabla g_{0})+\left(\eta_{\gamma}\nabla\log\frac{\eta_{\gamma}}{\pi}\right)\cdot\nabla g_{0}$
$-f_{\textrm{W}}(\eta_{\gamma})\left(g_{0}-\mathbb{E}_{\eta_{\gamma}}[g_{0}]\right)+\nabla\cdot\left(\eta_{\gamma}\nabla\log\frac{\eta_{\gamma}}{\pi}\right)+\eta_{\gamma}\mathbb{E}_{f_{\textrm{W}}(\eta_{\gamma})}[g_{0}]$
$=f_{\textrm{W}}(\eta_{\gamma})+\eta_{\gamma}\mathbb{E}_{f_{\textrm{W}}(\eta_{\gamma})}[g_{0}]+\nabla\cdot(\eta_{\gamma}\nabla g_{0})+\left(\eta_{\gamma}\nabla\log\frac{\eta_{\gamma}}{\pi}\right)\cdot\nabla g_{0}$
$=-\eta_{\gamma}\left|\nabla\log\frac{\eta_{\gamma}}{\pi}\right|^{2}-\eta_{\gamma}\int\log\frac{\eta_{\gamma}}{\pi}\nabla\cdot\left(\eta_{\gamma}\nabla\log\frac{\eta_{\gamma}}{\pi}\right)$
$=-\eta_{\gamma}\left(\left|\nabla\log\frac{\eta_{\gamma}}{\pi}\right|^{2}-\mathbb{E}_{\eta_{\gamma}}\left[\left|\nabla\log\frac{\eta_{\gamma}}{\pi}\right|^{2}\right]\right)$ (A.6)
@@
$=\frac{f_{\textrm{W}}(\eta_{\gamma})}{\eta_{\gamma}}-\left|\nabla\log\frac{\eta_{\gamma}}{\pi}\right|^{2}-\frac{1}{\eta_{\gamma}}\nabla\cdot\left(\eta_{\gamma}\nabla\log\frac{\eta_{\gamma}}{\pi}\right)$
@@
which has the same structure as (A.3) with $g_{0}$ now replaced by $g_{1}$. Nevertheless, we will calculate it to show this is the case. The Fréchet derivative of $f_{\textrm{W}},f_{\textrm{FR}}]$ is given by
@@
$=h(g_{1}-\mathbb{E}_{\eta_{\gamma}}[g_{1}])-\eta_{\gamma}\left(2\nabla\left(\frac{h}{\eta_{\gamma}}\right)\cdot\nabla\log\frac{\eta_{\gamma}}{\pi}-\int 2\nabla\left(\frac{h}{\eta_{\gamma}}\right)\cdot\nabla\log\frac{\eta_{\gamma}}{\pi}\eta_{\gamma}\right)-\eta_{\gamma}\int g_{1}h$
@@
where $g_{1}^{\prime}h:=-2\nabla\left(\frac{h}{\eta_{\gamma}}\right)\cdot\nabla\log\frac{\eta_{\gamma}}{\pi}$ is the Fréchet derivative of $g_{1}$ in the direction $h$. Notice this is the same structure as (A.5) but with $g_{0}$ replaced by $g_{1}$. Similarly, we can directly use (A.4) with $g_{1}$ in place of $g_{0}$ for the term $f_{\textrm{W}}^{\prime}[f_{\textrm{W}},f_{\textrm{FR}}](\eta_{\gamma})$ due to the Fisher-Rao structure of $f_{\textrm{W}},f_{\textrm{FR}}]$. This yields
@@
$=-\eta_{\gamma}(g_{1}^{\prime}f_{\textrm{W}}(\eta_{\gamma})-\mathbb{E}_{\eta_{\gamma}}[g_{1}^{\prime}f_{\textrm{W}}(\eta_{\gamma})])+\left(\eta_{\gamma}\nabla\log\frac{\eta_{\gamma}}{\pi}\right)\cdot\nabla g_{1}+\eta_{\gamma}\mathbb{E}_{f_{W}(\eta_{\gamma})}[g_{1}]+\nabla\cdot(\eta_{\gamma}\nabla g_{1})$
@@ Appendix B Splitting for multivariate Gaussian distributions
We recap here useful notation for this section:
  • $\mu_{0}(x)=\mathcal{N}(x;m_{0},C_{0})$ denotes the initial distribution and $\pi(x)=\mathcal{N}(x;m_{\pi},C_{\pi})$ the target distribution.
  • We denote by $\nu_{\gamma}$ the solution of the W-FR PDE in Proposition 2.2 for one time step $\gamma>0$.
  • We denote by $\eta_{\gamma}$ the solution of the FR-W PDE in Proposition 2.3 for one time step $\gamma>0$.
  • In the Gaussian case, $b_{\gamma},Q_{\gamma}$ denote mean and variance of $\nu_{\gamma}$.
  • In the Gaussian case, $a_{\gamma},P_{\gamma}$ denote mean and variance of $\eta_{\gamma}$.

### v2

@@ A.3 Auxiliary Lemmas
Define $f(\tau)=e^{(\gamma-\tau)f_{W}}\mathcal{G}^{1}\eta_{1}(x;\tau)$ and consider a Taylor expansion around point $\gamma$ of $f$, $f(\tau)=\sum_{k=0}^{\infty}\frac{f^{(k)}(\gamma)}{k!}(\tau-\gamma)^{k}$. Let us compute the derivatives of $f$. With a slight abuse of notation, in the remainder of the proof we use $\eta_{s}$ to denote $\eta_{1}(x;s)$ to emphasise dependence on $s$. Applying the chain rule we have
@@
$\partial_{s}f_{\textrm{W}}(f_{\textrm{FR}}(\eta_{s}))=f_{\textrm{W}}^{\prime}(f_{\textrm{FR}}(\eta_{s}))\left[\partial_{s}f_{\textrm{FR}}(\eta_{s})\right]=f_{\textrm{W}}(f^{\prime}_{\textrm{FR}}(\eta_{s})[\partial_{s}\eta_{s}])=f_{\textrm{W}}(f^{\prime}_{\textrm{FR}}(\eta_{s})[f_{\textrm{W}}(\eta_{s})])$ (1.8)
@@
$\partial_{s}f_{\textrm{FR}}^{\prime}(\eta_{s})[f_{\textrm{W}}(\eta_{s})]$ $=f_{\textrm{FR}}^{\prime\prime}(\eta_{s})(\partial_{s}\eta_{s},f_{\textrm{W}}(\eta_{s}))+f_{\textrm{FR}}^{\prime}(\eta_{s})[\partial_{s}f_{\textrm{W}}(\eta_{s})]$ (1.9)
@@
Combining (1.8)–(1.9) it the immediately follows that
@@
$\partial_{s}\mathcal{G}^{1}\eta_{s}-f_{\textrm{W}}(\mathcal{G}^{1})(\eta_{s})$ $=-[f_{\textrm{W}},[f_{\textrm{W}},f_{\textrm{FR}}]](S_{\textrm{W}}(s,\hat{\eta}_{1}(x;\gamma)))=-\mathcal{G}^{2}\eta_{s},$
@@
using the chain rule for operators. Following the same steps as before we have $\partial_{s}\mathcal{G}^{2}\eta_{s}-f_{\textrm{W}}(\mathcal{G}^{2}\eta_{s})=-\mathcal{G}^{3}\eta_{s}$ and thus $f^{(2)}(s)=e^{(\gamma-s)f_{W}}\mathcal{G}^{3}\eta_{s}$. A recursive argument then gives
@@
$=\gamma\mathcal{G}^{1}\eta_{\gamma}+\sum_{k=1}^{\infty}\frac{\mathcal{G}^{k+1}\eta_{\gamma}}{k!}\frac{\gamma^{k+1}}{k+1}=\sum_{k=1}^{\infty}\frac{\gamma^{k}}{k!}\mathcal{G}^{k}(\eta_{\gamma}).$
@@
Let Assumption 1 hold. Recall $\eta_{1}(x;\gamma)$ denotes the solution of (2.12) over a single step of size $\gamma$. Then the $k$th commutator $\mathcal{G}^{k}\eta_{1}:=[f_{\textrm{W}},\mathcal{G}^{k-1}\eta_{1}]$ for $k=1,2,3,\dots$ with $\mathcal{G}^{0}:=f_{\textrm{FR}}$ and $\mathcal{G}^{1}$ as defined in (1.5) has a Fisher-Rao structure,
$\mathcal{G}^{k}\eta_{1}=\eta_{1}\left(g_{k}(\eta_{1})-\mathbb{E}_{\eta_{1}}[g_{k}]\right)$
where $g_{k-1}^{\prime}f_{\textrm{W}}(\eta)$ denotes the directional derivative of $g_{k-1}$ in the direction $f_{\textrm{W}}(\eta)$ and
$g_{k}(\eta_{1})$ $=-g_{k-1}^{\prime}f_{\textrm{W}}(\eta_{1})+\nabla\log\frac{\eta_{1}}{\pi}\cdot\nabla g_{k-1}(\eta_{1})+\frac{1}{\eta_{1}}\nabla\cdot(\eta_{1}\nabla g_{k-1}(\eta_{1})),\quad\text{for}\enskip k=1,2,3,\dots$
$g_{0}(\eta_{1})$ $=-\log\frac{\eta_{1}}{\pi}$
@@
Once again, with a slight abuse of notation we use $\eta_{\gamma}$ to denote $\eta_{1}(x;\gamma)$ to emphasise the dependence on the step size $\gamma$. We will proceed inductively, starting with $\mathcal{G}^{1}$,
$[f_{\textrm{W}},f_{\textrm{FR}}]\eta_{\gamma}$ $=f_{\textrm{W}}^{\prime}f_{\textrm{FR}}(\eta_{\gamma})-f_{\textrm{FR}}^{\prime}f_{\textrm{W}}(\eta_{\gamma}).$ (1.10)
Since $f_{\textrm{W}}$ is a bounded linear operator under Assumption (a), its directional derivative is itself and
$f_{\textrm{W}}^{\prime}f_{\textrm{FR}}(\eta_{\gamma})$ $=\nabla\cdot\left(\nabla\eta_{\gamma}(g_{0}-\mathbb{E}_{\eta_{\gamma}}[g_{0}])\right)-\nabla\cdot\left(\eta_{\gamma}(g_{0}-\mathbb{E}_{\eta_{\gamma}}[g_{0}])\nabla\log\pi\right)$ (1.11)
@@
The directional derivative of $f_{\textrm{FR}}$ is given by
$\frac{d}{d\epsilon}\left.f_{\textrm{FR}}(\eta_{\gamma}+\epsilon h)\right|_{\epsilon=0}$ $=\frac{d}{d\epsilon}\left.-(\eta_{\gamma}+\epsilon h)\log\frac{\eta_{\gamma}+\epsilon h}{\pi}+(\eta_{\gamma}+\epsilon h)\int\log\frac{\eta_{\gamma}+\epsilon h}{\pi}(\eta_{\gamma}+\epsilon h)\right|_{\epsilon=0}$ (1.12)
@@
Then since $g_{0}^{\prime}h=-\frac{h}{\eta_{\gamma}}$,
@@
$[f_{\textrm{W}},f_{\textrm{FR}}]\eta_{\gamma}$ $=f_{\textrm{W}}(\eta_{\gamma})+\eta_{\gamma}\mathbb{E}_{f_{\textrm{W}}(\eta_{\gamma})}[g_{0}]+\nabla\cdot(\eta_{\gamma}\nabla g_{0})+\left(\eta_{\gamma}\nabla\log\frac{\eta_{\gamma}}{\pi}\right)\cdot\nabla g_{0}$
$=-\eta_{\gamma}\left(\left|\nabla\log\frac{\eta_{\gamma}}{\pi}\right|^{2}-\mathbb{E}_{\eta_{\gamma}}\left[\left|\nabla\log\frac{\eta_{\gamma}}{\pi}\right|^{2}\right]\right)$ (1.13)
@@
which has the same structure as (1.10) with $g_{0}$ now replaced by $g_{1}$. Nevertheless, we will calculate it to show this is the case. The directional derivative of $f_{\textrm{W}},f_{\textrm{FR}}]$ is given by
@@
where $g_{1}^{\prime}h:=-2\nabla\left(\frac{h}{\eta_{\gamma}}\right)\cdot\nabla\log\frac{\eta_{\gamma}}{\pi}$ is the directional derivative of $g_{1}$ in the direction $h$. Notice this is the same structure as (1.12) but with $g_{0}$ replaced by $g_{1}$. Similarly, we can directly use (1.11) with $g_{1}$ in place of $g_{0}$ for the term $f_{\textrm{W}}^{\prime}[f_{\textrm{W}},f_{\textrm{FR}}](\eta_{\gamma})$ due to the Fisher-Rao structure of $f_{\textrm{W}},f_{\textrm{FR}}]$. This yields
@@
$=-\eta_{\gamma}(g_{1}^{\prime}f_{\textrm{W}}(\eta_{\gamma})-\mathbb{E}_{\eta_{\gamma}}[g_{1}^{\prime}f_{\textrm{W}}(\eta_{\gamma})])+\left(\eta_{\gamma}\nabla\log\frac{\eta_{\gamma}}{\pi}\right)\cdot\nabla g_{1}+\eta_{\gamma}\mathbb{E}_{f_{W}(\eta_{\gamma})}[g_{1}]+\nabla\cdot(\eta_{\gamma}\nabla g_{1}).$
@@ Appendix B Splitting for multivariate Gaussian distributions
In this section, $\mu_{0}(x)=\mathcal{N}(x;m_{0},C_{0})$ and $\pi(x)=\mathcal{N}(x;m_{\pi},C_{\pi})$. Also, $b_{1},Q_{1}$ denote mean and covariance of $\nu_{1}$, the solution of the W-FR scheme after one step and $a_{1},P_{1}$ denote mean and covariance of $\eta_{1}$, the solution of the FR-W scheme after one step.